Multiply and then simplify completely: (V5 - 2)? Answer
Answer #1

If you have any queries please let me know in comments and rate it.

(15 - 2)² (a - b) = a +6²-2ab. a=rs, b=52 (11) ² (343) 2axus xv= = 5+ 2- 2550 (55-1E)² -7-2010